OpenStudy (anonymous):

A truck accelerates from rest to a speed of 17 m/s in 9.1 s. If the tires have a radius of 33 cm, how many revolutions do the tires make during this time

OpenStudy (shamim):

u need to use formula as below v=u+at v^2=u^2+2as circumferance=2*pi*r revolution no.=s/2*pi*r

OpenStudy (anonymous):

i got this question from the help i got when i last posted it.

OpenStudy (shamim):

ok

OpenStudy (anonymous):

can you define your variables??

OpenStudy (shamim):

ya v= final velocity u= initial velocity s=distance travelled a=acceleration r=radius of tyre t= time

OpenStudy (anonymous):

thanks I got it right this time.

OpenStudy (shamim):

welcome

OpenStudy (shamim):

welcome